# A new economic and financial theory of money

### v6

@@ VII Solution of the Hamilton-Jacobi-Bellman equation
The motion, at this point, has been transformed into the space of all possible local complex spectrums $\left|\beta(z)\right>$ (that is, all possible solutions to the renormalization group equations). This expansion or domain can be interpreted several different ways in addition to as solutions to the renormalization group equations. It can be interpreted as Heisenberg’s S-matrix expansion, as the $m$-body scattering cross sections, as the $m$-body Green’s functions, and the Mayer Cluster Expansion. This is why it has been called the Heisenberg Scattering Transformation. It captures the $m$-body correlation structure of the collective motion. This correlation synchronizes the collective motion so that the collective acts as one. Motion of the system (or variation in $H(z)$) will be confined to a $n$-dimensional complex linear subspace $\mathbb{C}^{n}$ of the individual motion with basis vectors $\left|\beta_{i}(z)\right>$ easily identified by a Principal Components Analysis (PCA). The $\left|\beta_{i}(z)\right>$ are the solutions to the renormalization group equations for the field theory with Lagrangian functional $L[f(x),\dot{f}(x)]$. They are also the complex curvature or S-matrix $S_{m}$ given in Eq.  91. The motion is projected onto this complex linear sub-space with complex coordinates $\beta=(\{\beta_{i}\})$.
The connection of the HST detailed in Eq. 101 to CNNs can be seen by identifying the iterative deep convolutional structure of the product, the nonlinear activation function $\text{i}\,\ln R_{0}$, and the pooling operation $\phi_{px}$. The HST followed by the PCA projection to $\beta$ also can be interpreted as the Wigner-Weyl transformation done right (Case, 2008; Wigner, 1932; Weyl, 1950).
@@
Now, consider an external force exerted on a economic sub-system coming from an external economic heat bath of economic temperature $T_{e}\ll J_{0}\,\omega_{0}$, where $J_{0}$ is the economic activity of the economic sub-system and $\omega_{0}$ is the billing frequency of the economic sub-system. This leads to a diffusion in $P$ governed by the Black-Scholes or Fokker-Plank equation shown in Eq. 117, where $\nu_{c}$ is the rate at which steps are taken in the economic sub-system of size $\sigma_{c}$ due to the external interaction. The equilibrium distribution will be
@@
with the caveat that $P^{*}_{s}$ is a stable or stabilized equilibrium. The status quo is that there is no stabilizing force $F_{sf}(P)$ or $F_{sp}(\tau)$ of unstable equilibriums or control force $F_{c}(q)$ to optimize the desired reward $R(q)$ so the economic sub-system relaxes to undesired stable equilibriums (local minimums) with $T_{e}$ giving the natural prime interest rate $\nu_{\text{prime}}=\sqrt{T_{e}/E_{0}}\,\omega_{0}$.
@@ IX Non-conservative HJB Equation
It is very common to embed an external non-conservative force $F_{\text{ext}}(\omega_{Q},Q)$ in the sub-system dynamics instead of applying the external non-conservative force to the sub-system. In many cases the external force is resistive, that is a thermal adiabatic force $F_{\text{ext}}(\omega_{Q},Q)=-\epsilon_{P}\,\omega_{Q}$. The effect on the dynamics of embedding the external force rather than applying the external force is dramatic. Applying an external resistive force to a conservative sub-system simply relaxes the sub-system to the equilibriums of the conservative dynamics where the external system no longer has an influence on the dynamics and the sub-system follows the conservative dynamics. The influence of the external resistive force, in fact any external force, can be both mitigated and controlled. Embedding an external force fundamentally degrades the geometry, and eventually changes and destroys the topology of the dynamics. If the external force is resistive, the topology is regularized. If the regularization is small $\nu\lesssim\omega_{0}$, then the topology is preserved, but the energy of the equilibriums $E_{0}$ are reduced as shown in Fig. 14. If the regularization is large $\nu>\nu_{\text{cr}}\gg\omega_{0}$, the topology is reduced to one with a single homology class of a point so that trajectories inexorably spiral to the point at zero with zero energy.
@@
To understand what is happening to the dynamics when the external non-conservative force is applied, the HJB equation needs to be modified to include this force. The common way that this is done results in the following equation
@@
A constrained optimization of this $V[q(\tau)]$ gives the expression for NPV found in Eq. 1.
@@ X Relationship to Deep Reinforcement Learning
The art is in how to relax the value and policy to the optimum at $\theta=\theta^{*}$. The most successful strategy has been Q-learning. This strategy is based on the following equation obtained by substituting the expression for $\widetilde{V}(s;\theta)$ in Eq. 133 into the definition of $\widetilde{Q}(s,a;\theta)$ in Eq. 132
@@
Before the differences between this traditional approach and the methodology proposed in this paper for optimal control can be understood, the relationship of the discrete Bellman equation shown in Eq. 131 to the continuous HJB shown in Eq. 124 needs to be established. Start with the Bellman equation, Eq. 131, and rearrange terms to give
@@
with a high performance root solver, rather than the much slower gradient decent of Eq. 148. Then, the optimal value function
@@
Once these surrogates are formed, the equilibrium points $P^{*}$ can be efficiently found in 10’s of seconds. Not only are the stable equilibriums (that is local minimums, o-points, or attractors) found, but the more desirable unstable equilibriums (that is local maximums, x-points, saddle points, or semi-attractors) are also found. These unstable equilibriums can then be stabilized and protected from external forces by the use of Eq. 100 or Eq. 99. The reward function can also be modified as shown in Eq. 88. DRL uses a much slower gradient decent that needs to calculate two time-consuming minimums or maximums for each sample.
@@
DRL finds the optimum policy $\pi^{*}(q)$, not the control forces, $F_{c}(q)$, $F_{sf}(P)$, and $F_{sp}(\tau)$. The optimum policy is neither single valued nor existing for all $q$. This manifests itself in the DRL algorithm as a lack of convergence, lack of stationarity, and lack of IID (Independent and Identical Distribution). DQN resolves these problems by an opaque and laborious procedure where
@@
the feedback stabilization force is learned, so that the force $\Delta p/\Delta\tau=-\omega_{sf}\Delta\pi(q,P)$ can be applied. It should be noted that the inputs to DQN are both $f(x)$ and $\pi(x)$, not just $f(x)$. This is done by inputting the field, effectively $q$, at different times so that $\dot{q}$, effectively the co-state $P$, can be estimated.
It is straight forward to construct training datasets using our methodology that sample phase space well and sufficiently explore off-policy. This can be done by integrating trajectories with a small external dissipative force. It is an exceptionally difficult task for DRL to keep most of the steps on-policy and simultaneously exploring off-policy, but not to far off-policy, for the determination of the adjoint motion. The challenge can be seen in the form of Eq. 136, where $\widetilde{Q}_{\text{target}}(s,a,;\theta)$ is neither stationary nor IID.
The diffusion from the external force is modelled correctly with the $F_{\text{ext}}(\omega_{Q},Q)$, retaining the geometry and topology of the conservative Hamiltonian system using our methodology. It is not embedded in the dynamics as shown in Eq. 124, first degrading the geometry, then destroying the topology of the conservative Hamiltonian system. The regularization, if it is small $\nu\lesssim\omega_{0}$, leads to a geometric degradation, significantly reducing $E_{0}$. If it is large $\nu>\nu_{\text{cr}}\gg\omega_{0}$, it leads to topological destruction and all trajectories spiraling to zero with zero energy. The ubiquitous use of the discount factor $\gamma$ in DRL, to improve convergence and regularize the solution (needed because of the incomplete reduction to the dynamical variables $q(\tau)$) has these unintended effects.
@@ XII Example of an electronic currency at work
In order to simplify this example we have separated the development into four phases: (1) startup, (2) development of the energy transition for the state of New Mexico, (3) development of the energy transition for the United States, and (4) steady state. For both Phase 2 and Phase 3, there are one NewCo and one BuyCo. In practice, there will be a portfolio of companies. It is assumed that \$20 million USD will be needed for Phase 1 and will be raised from the private markets at a valuation of \$60 million USD. It will be used to set up the EC\$, to implement the on-line real time financial accounting, to obtain the necessary US SEC approvals, to establish the legal contracts for operating the sub-economy, and for governmental and public relations. The beginning of Phase 2 will commence with an IPO to raise \$3 billion USD in cash reserves for the EC\$ at a New Energy Inc. valuation of \$15 billion USD. The BuyCo is an typical size existing producer of unconventional gas in the Permian basin, and the NewCo is a production plant of blue hydrogen located next to the production reservoirs of the BuyCo. Details of the transactions, including detailed balance sheets, can be found in App. A along with other details. An interesting thing to note on the summary of the balance sheets shown in Fig. 32 is that the entries on the balance sheet for New Energy Inc. are reversed like a central bank with the EC\$ loans appearing as assets and the EC\$ as an equity (liability).
@@ XIII Ubuntu business model
The Ubuntu business model recognizes that the economy is a collective of individuals, as such, there emerges a collective behavior or virtual individual that determines, that is controls, the collective behavior. This control emanates from a simply beautiful symmetry – the coordination of the economically interacting individuals creating an economic force, and the economic force then coordinating the economic trade of the individuals. Another way of looking at this is that the actions of the virtual individual, who follows a geodesic motion, are approximated by genAI. The genAI then generates and multiply reflects the actions of the virtual individual through a Hall of Mirrors to obtain, that is forecast, the economic evolution of the collective. This is demonstrated by “The Mirror Maze” scene from the 1928 Charlie Chaplin movie “The Circus” (Chaplin, 1928) and shown in Fig. 22.
@@
The Ubuntu way of “mutuality” is based on the concept of sharing, putting toys in a community toy chest so that there are more toys from which to choose as shown in Fig. 23. The community software chest is GitHub. The community capital chest is Martin Luther’s Community Chest of the Protestant Reformation (Luther, 1523). The electronic transaction equity version of community capital chest is Ubuntu Financing. This is related to the Economics of Mutuality (Mayer and Roche, 2021), Modern Monetary Theory (Kelton, 2020), and Special Purpose Acquisition Companies (SPACs).
Figure 23: The good of sharing. (a) Children putting toys in a community toy chest so that there are more toys from which to choose, leading to happiness. (b) Children being greedy, keeping their toys to themselves, leading to sadness.
@@ XIV Government and taxation
For games and economic systems, the system dynamics are determined and/or modified by rules and regulations. These can either be by modification of the reward or conservative potential economic energy $R(q)$ or by modification of the force or non-conservative external interaction $f(q,\dot{q})$. Even though the force is non-conservative, the system will remain conservative. For this reason, it is very important that we digress at this time and closely examine what rules, regulations, and taxes will keep the system conservative, and which ones will not. This also informs some of the freedoms that need to be protected to keep the system conservative.
@@
so that $f$ can be time dependant, but $R$ can not and still remain a conservative system. So, to remain conservative, the reward or potential must be a function of $q$ only $R(q)$. It can not be a function of either $p$ or $\tau$. There is no restriction on the form of $f(p,q,\tau)$. The reward can be identified by the energy $E$ when $p=a=0$, that is no action is being taken. This energy or reward when there is no action $E(p=0,q)$ can not be a function of time. The restriction of being a force on the system is that $\Delta R=\dot{q}\,f(p,q,\tau)\,d\tau$. In other words, the reward is proportional to the rate of change in the state – a tax on activity.
@@
The approach to management should be that of a private corporation with government appointed representatives forming the board of directors. There should be separate corporations for the federal, state and local governments. As usual, it is the board of directors that approves the fiscal policy (that is budget) of the corporation proposed by the executives of the corporation, Government Inc. The budget could be subject to the approval of the respective elected legislative branch of government, and the key chief executives could be the elected members of the executive branch of government. There should be an electronic currency for each corporation. The monetary policy should be coordinated with the fiscal policy and implemented in the manner proposed by this paper.
@@ XV Freedom of labor, intellectual property, and law
We start by discussing the concept of “Communism & Freedom”. Milton Friedman went out of his way in the early 1960’s to cast capitalism as being superior to communism, but he made the point that it was essential that there be freedom so that the private sector could zealously pursue its maximization of profit. If it was profitable, then it was in the social good. This is fundamentally flawed. Greed is not good, it is evil. What is profitable is exploiting society. The resistivity that the pursuit of profitability imposes on society, leads to oppression that opposes freedom (Glinsky, 2024b). Friedman’s ideas formed the basis of Reagan and Thatcher conservatism of the 1980’s.

### v7

@@ VII Solution of the Hamilton-Jacobi-Bellman equation
The motion, at this point, has been transformed into the space of all possible local complex spectrums $\left|\beta(z)\right>$ (that is, all possible solutions to the renormalization group equations). This expansion or domain can be interpreted several different ways in addition to as solutions to the renormalization group equations. It can be interpreted as Heisenberg’s S-matrix expansion, as the $m$-body scattering cross sections, as the $m$-body Green’s functions, and the Mayer Cluster Expansion. This is why it has been called the Heisenberg Scattering Transformation. It captures the $m$-body correlation structure of the collective motion. This correlation synchronizes the collective motion so that the collective acts as one. Motion of the system (or variation in $H(z)$) will be confined to a $n$-dimensional complex linear subspace $\mathbb{C}^{n}$ of the individual motion with basis vectors $\left|\beta_{i}(z)\right>$ easily identified by a Principal Components Analysis (PCA). The $\left|\beta_{i}(z)\right>$ are the solutions to the renormalization group equations for the field theory with Lagrangian functional $L[f(x),\dot{f}(x)]$. They are also the complex curvature or S-matrix $S_{m}$ given in Eq. (91). The motion is projected onto this complex linear sub-space with complex coordinates $\beta=(\{\beta_{i}\})$.
The connection of the HST detailed in Eq. (101) to CNNs can be seen by identifying the iterative deep convolutional structure of the product, the nonlinear activation function $\text{i}\,\ln R_{0}$, and the pooling operation $\phi_{px}$. The HST followed by the PCA projection to $\beta$ also can be interpreted as the Wigner-Weyl transformation done right (Case, 2008; Wigner, 1932; Weyl, 1950).
@@
Now, consider an external force exerted on a economic sub-system coming from an external economic heat bath of economic temperature $T_{e}\ll J_{0}\,\omega_{0}$, where $J_{0}$ is the economic activity of the economic sub-system and $\omega_{0}$ is the billing frequency of the economic sub-system. This leads to a diffusion in $P$ governed by the Black-Scholes or Fokker-Plank equation shown in Eq. (117), where $\nu_{c}$ is the rate at which steps are taken in the economic sub-system of size $\sigma_{c}$ due to the external interaction. The equilibrium distribution will be
@@
with the caveat that $P^{*}_{s}$ is a stable or stabilized equilibrium. The status quo is that there is no stabilizing force $F_{\text{sf}}(P)$ or $F_{\text{sp}}(\tau)$ of unstable equilibriums or control force $F_{c}(q)$ to optimize the desired reward $R(q)$ so the economic sub-system relaxes to undesired stable equilibriums (local minimums) with $T_{e}$ giving the natural prime interest rate $\nu_{\text{prime}}=\sqrt{T_{e}/E_{0}}\,\omega_{0}$.
@@ IX Non-conservative HJB Equation
It is very common to embed a non-conservative force in the sub-system dynamics instead of applying an external force to the sub-system. The effect on the dynamics of embedding a non-conservative force rather than applying an external force is dramatic. Applying an external resistive force to a conservative sub-system simply relaxes the sub-system to the equilibriums of the conservative dynamics where the external system no longer has an influence on the dynamics and the sub-system follows the conservative dynamics. The influence of the external resistive force, in fact any external force, can be both mitigated and controlled. Embedding a non-conservative force fundamentally degrades the geometry, and eventually changes and destroys the topology of the dynamics. If the embedded non-conservative force is resistive, the topology is regularized. If the regularization is small $\nu\lesssim\omega_{0}$, then the topology is preserved, but the energy of the equilibriums $E_{0}$ are reduced as shown in Fig. 14. If the regularization is large $\nu>\nu_{\text{cr}}\gg\omega_{0}$, the topology is reduced to one with a single homology class of a point so that trajectories inexorably spiral to the point at zero with zero energy.
@@
To understand what is happening to the dynamics when the non-conservative force is embedded, the HJB equation needs to be modified to include this force. The common way that this is done results in the following equation
@@
A constrained optimization of this $V[q(\tau)]$ gives the expression for NPV found in Eq. (1).
@@ X Relationship to Deep Reinforcement Learning
The art is in how to relax the value and policy to the optimum at $\theta=\theta^{*}$. The most successful strategy has been Q-learning. This strategy is based on the following equation obtained by substituting the expression for $\widetilde{V}(s;\theta)$ in Eq. (133) into the definition of $\widetilde{Q}(s,a;\theta)$ in Eq. (132)
@@
Before the differences between this traditional approach and the methodology proposed in this paper for optimal control can be understood, the relationship of the discrete Bellman equation shown in Eq. (131) to the continuous HJB shown in Eq. (124) needs to be established. Start with the Bellman equation, Eq. (131), and rearrange terms to give
@@
with a high performance root solver, rather than the much slower gradient decent of Eq. (148). Then, the optimal value function
@@
Once these surrogates are formed, the equilibrium points $P^{*}$ can be efficiently found in 10’s of seconds. Not only are the stable equilibriums (that is local minimums, o-points, or attractors) found, but the more desirable unstable equilibriums (that is local maximums, x-points, saddle points, or semi-attractors) are also found. These unstable equilibriums can then be stabilized and protected from external forces by the use of Eq. (100) or Eq. (99). The reward function can also be modified as shown in Eq. (88). DRL uses a much slower gradient decent that needs to calculate two time-consuming minimums or maximums for each sample.
@@
DRL finds the optimum policy $\pi^{*}(q)$, not the control forces, $F_{c}(q)$, $F_{\text{sf}}(P)$, and $F_{\text{sp}}(\tau)$. The optimum policy is neither single valued nor existing for all $q$. This manifests itself in the DRL algorithm as a lack of convergence, lack of stationarity, and lack of IID (Independent and Identical Distribution). DQN resolves these problems by an opaque and laborious procedure where
@@
the feedback stabilization force is learned, so that the force $\Delta p/\Delta\tau=-\omega_{\text{sf}}\Delta\pi(q,P)$ can be applied. It should be noted that the inputs to DQN are both $f(x)$ and $\pi(x)$, not just $f(x)$. This is done by inputting the field, effectively $q$, at different times so that $\dot{q}$, effectively the co-state $P$, can be estimated.
It is straight forward to construct training datasets using our methodology that sample phase space well and sufficiently explore off-policy. This can be done by integrating trajectories with a small external dissipative force. It is an exceptionally difficult task for DRL to keep most of the steps on-policy and simultaneously exploring off-policy, but not to far off-policy, for the determination of the adjoint motion. The challenge can be seen in the form of Eq. (136), where $\widetilde{Q}_{\text{target}}(s,a,;\theta)$ is neither stationary nor IID.
The diffusion from the external force is modelled correctly with the $F_{\text{ext}}(\omega_{Q},Q)$, retaining the geometry and topology of the conservative Hamiltonian system using our methodology. It is not embedded in the dynamics as shown in Eq. (124), first degrading the geometry, then destroying the topology of the conservative Hamiltonian system. The regularization, if it is small $\nu\lesssim\omega_{0}$, leads to a geometric degradation, significantly reducing $E_{0}$. If it is large $\nu>\nu_{\text{cr}}\gg\omega_{0}$, it leads to topological destruction and all trajectories spiraling to zero with zero energy. The ubiquitous use of the discount factor $\gamma$ in DRL, to improve convergence and regularize the solution (needed because of the incomplete reduction to the dynamical variables $q(\tau)$) has these unintended effects.
@@ XII Example of an electronic currency at work
In order to simplify this example we have separated the development into four phases: (1) startup, (2) development of the energy transition for the state of New Mexico, (3) development of the energy transition for the United States, and (4) steady state. For both Phase 2 and Phase 3, there are one NewCo and one BuyCo. In practice, there will be a portfolio of companies. It is assumed that \$20 million USD will be needed for Phase 1 and will be raised from the private markets at a valuation of \$60 million USD. It will be used to set up the EC\$, to implement the on-line real time financial accounting, to obtain the necessary US SEC approvals, to establish the legal contracts for operating the sub-economy, and for governmental and public relations. The beginning of Phase 2 will commence with an IPO to raise \$3 billion USD in cash reserves for the EC\$ at a New Energy Inc. valuation of \$15 billion USD. The BuyCo is a typical size existing producer of unconventional gas in the Permian basin, and the NewCo is a production plant of blue hydrogen located next to the production reservoirs of the BuyCo. Details of the transactions, including detailed balance sheets, can be found in App. A along with other details. An interesting thing to note on the summary of the balance sheets shown in Fig. 32 is that the entries on the balance sheet for New Energy Inc. are reversed like a central bank with the EC\$ loans appearing as assets and the EC\$ as an equity (liability).
@@ XIII Ubuntu business model
The Ubuntu business model recognizes that the economy is a collective of individuals, as such, there emerges a collective behavior or virtual individual that determines, that is controls, the collective behavior (Glinsky, 2024d). This control emanates from a simply beautiful symmetry – the coordination of the economically interacting individuals creating an economic force, and the economic force then coordinating the economic trade of the individuals. Another way of looking at this is that the actions of the virtual individual, who follows a geodesic motion, are approximated by genAI. The genAI then generates and multiply reflects the actions of the virtual individual through a Hall of Mirrors to obtain, that is forecast, the economic evolution of the collective. This is demonstrated by “The Mirror Maze” scene from the 1928 Charlie Chaplin movie “The Circus” (Chaplin, 1928) and shown in Fig. 22.
@@
The Ubuntu way of “mutuality” is based on the concept of sharing, putting toys in a community toy chest so that there are more toys from which to choose as shown in Fig. 23. The community software chest is GitHub. The community capital chest is Martin Luther’s Community Chest of the Protestant Reformation (Luther, 1523). The electronic transaction equity version of the community capital chest is Ubuntu Financing. This is related to the Economics of Mutuality (Mayer and Roche, 2021), Modern Monetary Theory (Kelton, 2020), Marxian Economics (Wolff, 1987; Desai, 2006; mar, 2024; Douglas, 1920), Complexity Economics (Farmer, 2024), and Special Purpose Acquisition Companies (SPACs) (spa, 2024; Ren, 2021).
Figure 23: The good of sharing. (a) Children putting toys in a community toy chest so that there are more toys from which to choose, leading to happiness. (b) Children being greedy, keeping their toys to themselves, leading to sadness. Images generated by OpenAI’s DALL-E.
@@ XIV Government and taxation
For games and economic systems, the system dynamics are determined and/or modified by rules and regulations. These can either be by modification of the reward or conservative potential economic energy $R(q)$, or by application of an external force $f(q,\dot{q})$. Even though the external force does work on the system and changes its energy, the system will remain conservative. For this reason, it is very important that we digress at this time and closely examine what rules, regulations, and taxes will keep the system conservative, and which ones will not. This also informs some of the freedoms that need to be protected to keep the system conservative.
@@
so that $f$ can be time dependant, but $R$ can not and still remain a conservative system. So, to remain conservative, the reward or potential must be a function of $q$ only, that is $R(q)$. It can not be a function of either $p$ or $\tau$. There is no restriction on the form of $f(p,q,\tau)$. The reward can be identified by the energy $E$ when $p=a=0$, that is no action is being taken. This energy or reward when there is no action $E(p=0,q)$ can not be a function of time. The restriction of being a force on the system is that $\Delta R=\dot{q}\,f(p,q,\tau)\,d\tau$. In other words, the reward is proportional to the rate of change in the state – a tax on activity.
@@
The approach to management of government actions should be that of a private corporation with government appointed representatives forming the board of directors. There should be separate corporations for the federal, state and local governments. As usual, it is the board of directors that approves the fiscal policy (that is budget) of the corporation proposed by the executives of the corporation, Government Inc. The budget could be subject to the approval of the respective elected legislative branch of government, and the key chief executives could be the elected members of the executive branch of government. There should be an electronic currency for each corporation. The monetary policy should be coordinated with the fiscal policy and implemented in the manner proposed by this paper.
@@ XV Freedom of labor, intellectual property, and law
We start by discussing the concept of “Communism & Freedom”. Milton Friedman went out of his way in the early 1960’s to cast capitalism as being superior to communism, but he made the point that it was essential that there be freedom so that the private sector could zealously pursue its maximization of profit. If it was profitable, then it was in the social good. This is fundamentally flawed. Greed is not good, it is evil. What is profitable is exploiting society. The resistivity that the pursuit of profitability imposes on society, leads to oppression that opposes freedom (Glinsky, 2024d). Friedman’s ideas formed the basis of Reagan and Thatcher conservatism of the 1980’s.